# Stationary solutions to stochastic 3D Euler equations in Hölder space

### v1

@@ 2.2. Probabilistic elements
For a given probability measure $\mathbf{P}$ we denote by $\mathbf{E}$ the expectation under $\mathbf{P}$. Regarding the driving noise, we assume that $B$ is $\mathbb{R}^{3}$-valued two-sided $GG^{*}$-Wiener process with zero spatial mean and zero divergence, defined on some probability space $(\Omega,\mathcal{F},\mathbf{P})$ and $G$ is a Hilbert–Schmidt operator from $U$ to $L_{\sigma}^{2}$ for some Hilbert space $U$. For $p\in[1,\infty)$, $\delta\in(0,\frac{1}{2})$ and some $\kappa>0$, we denote
$|\!|\!|u|\!|\!|^{p}_{C^{0},p}:=\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{s\in[t,t+1]}\|u(s)\|^{p}_{C^{0}}\right],\quad|\!|\!|u|\!|\!|^{p}_{C_{t}^{\frac{1}{2}-\delta}C^{\kappa},p}:=\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|u\|^{p}_{C_{t}^{\frac{1}{2}-\delta}C^{\kappa}}\right].$
These norms define function spaces of random variables on $\Omega$ taking values in $C({\mathbb{R}}\times\mathbb{T}^{3})$ and $C^{\frac{1}{2}-\delta}({\mathbb{R}},C^{\kappa})$, respectively, with bounds in $L^{p}(\Omega;{C^{0}(I\times\mathbb{T}^{3}))}$ and $L^{p}(\Omega;{C^{\frac{1}{2}-\delta}(I,C^{\kappa})})$ for bounded interval $I\subset{\mathbb{R}}$. Furthermore, the bounds only depend on the length of the interval $I$, not on its location within ${\mathbb{R}}$. In the sequel, we simply say that $u$ has a uniform moment of order $p$ locally in $C^{0}({\mathbb{R}}\times\mathbb{T}^{3})$ provided $|\!|\!|u|\!|\!|^{p}_{C^{0},p}<\infty$. Similarly, we define the corresponding norms with $C^{0}$ replaced by $C^{\kappa}$, $C^{1}$, $L^{1}$ and $H^{\frac{3}{2}+\kappa}$.
@@ 3. Stochastic convex integration with pathwise estimates and results
In [CDZ22] and [HZZ22b], they presented a stochastic convex integration method constructing directly solutions on the whole time line $\mathbb{R}$ by introducing expectations to the iterative estimates in convex integration. In this section, we integrate their methods with pathwise estimates in the context of the stochastic Euler equations. Our goal is to construct a Hölder continuous Euler flow, mirroring the deterministic case as presented in [BV19], where the transport equations (see (4.7) below) are used to improve regularity. This method necessitates pathwise positive lower bounds for the solutions of the transport equations. As a result, moment bounds during the inductive iteration become insufficient. Hence, we introduce additional pathwise estimates in convex integration schemes. More precisely, our idea employs the cutoff method to control the growth of the noise, enabling the introduction of pathwise estimates during the inductive iteration. This is combined with moment bounds to derive desired uniform estimates. Compared to only using moment bounds in the inductive iteration in [CDZ22] and [HZZ22b], the advantage of this approach is that it only requires lower moments of the solutions, and avoids the requirements for higher moments as in [HZZ22b]. This can simplify the computations and enhance the regularity of the solutions. At the first step, we decompose stochastic Euler system (1.1) into two parts, one is linear and involves the noise, whereas the second one is a random PDE. More precisely, we consider the stochastic linear equation
@@
where $B$ is $\mathbb{R}^{3}$-valued two-sided trace-class $GG^{*}$-Wiener process with spatial zero mean and divergence free (see e.g. [LR15, Section 2.1]), and $v$ solves the nonlinear equation
@@
Here, $z$ is divergence free by the assumptions on the noise $B$ and by $P$ we denote the pressure term associated to $v$. We include the linear zero order term $z$ on the left hand side of (3.1) in order to obtain a stabilization of the equation needed for the necessary global in time estimates. It will be seen in the course of the construction that the corresponding counter term in (3.2) will not cause any difficulties. Using the factorization method, it is standard to derive regularity of the stochastic convolution $z$ on a given stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ canonical filtration given in [LR15, Section 2.1]. In particular, the following result follows from [DPZ92, Theorem 5.16] together with the Kolmogorov continuity criterion.
@@
We recall that the unique stationary solution to (3.1) has the explicit form $z(t)=\int_{-\infty}^{t}e^{s-t}{\mathord{{\rm d}}}B_{s}$. By [LR15, Section 2.1], the Wiener process $B$ can be written as $B=\sum_{k\in{\mathbb{N}}}\sqrt{c_{k}}\beta_{k}e_{k}$ for an orthonormal basis $\{e_{k}\}_{k\in{\mathbb{N}}}$ of $L^{2}_{\sigma}$ consisting of eigenvectors of $GG^{*}$ with corresponding eigenvalues $c_{k}$, and the coefficients satisfy $\sum_{k\in{\mathbb{N}}}c_{k}<\infty$, where $\{\beta_{k}\}_{k\in{\mathbb{N}}}$ are a sequence of mutually independent standard two-sided real-valued Brownian motions. Then it holds for $t\geqslant s$,
@@
The result then follows from Kolmogorov’s continuity criterion and fundamental Sobolev embedding $H^{\frac{3}{2}+\kappa}\hookrightarrow C^{\kappa}$, i.e. for any $\delta\in(0,\frac{1}{2})$
@@ 3.1. Stochastic convex integration set-up and iterative proposition
Let us now explain how the convex integration iteration is set up. We consider an increasing sequence $\{\lambda_{q}\}_{q\in\mathbb{N}_{0}}$ which diverges to $\infty$, and a bounded sequence $\{\delta_{q}\}_{q\in\mathbb{N}}$ which decreases to 00. We choose $a,b\in\mathbb{N}$ and $\beta\in(0,1)$ and let
@@
Here $\beta$ will be chosen sufficiently small, $a$ will be chosen sufficiently large, and more details on the choice of these parameters will be given below in the course of the construction. The iteration is indexed by a parameter $q\in\mathbb{N}_{0}$. At each step $q$, a pair $(v_{q},\mathring{R}_{q})$ is constructed to solve the following system
@@
where $\chi_{q}$ is a non-increasing smooth function such that
$\chi_{q}(x)=\begin{cases}1,&x\in[0,\lambda_{q+2}^{\gamma}],\\ 0,&x\in(\lambda_{q+3}^{\frac{1}{4}\gamma},\infty),\end{cases}$
with derivative bounded by $1$, which requires
$\frac{1}{\lambda_{q+3}^{\gamma/4}-\lambda_{q+2}^{\gamma}}\leqslant 1.$ (3.6)
And $\tilde{z}_{q}=\mathbb{P}_{\leqslant f(q)}z$ with $f(q)=\lambda_{q+1}^{\gamma/8}$. Here $\mathbb{P}_{\leqslant f(q)}$ is the Fourier multiplier operator, which projects a function onto its Fourier frequencies $\leqslant f(q)$ in absolute value, and $\gamma$ is another small parameter which will be fixed in Section 4.1. $\mathring{R}_{q}$ on the right hand of (3.4) is a $3\times 3$ matrix which is trace-free and we put the trace part into the pressure. Keeping (3.3) in mind, by the Sobolev embedding $\|f\|_{L^{\infty}}\lesssim\|f\|_{H^{\frac{3}{2}+\kappa}}$ for $\kappa>0$, we observe that for any $p\geqslant 2$ and $\delta\in(0,\frac{1}{2})$
@@
The above estimates (3.7) will be frequently used in the sequel. Under the above assumptions, our main iteration reads as follows, and the proof of this result is presented in Section 4 below.
@@
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$ and let $r>1$ be fixed. Given a smooth function $e:\mathbb{R}\to(0,\infty)$ such that $\bar{e}\geqslant e(t)\geqslant\underline{e}\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}L^{2}$ with $\|e^{\prime}\|_{C^{0}}\leqslant\tilde{e}$ for some constants $\overline{e},\underline{e},\tilde{e}>0$, where $L$ is from Proposition 3.2 and $c_{R}>0$ is a small universal constant (see Section 4.3). There exists a choice of parameters $a,b,\beta$ and $\gamma$ such that the following holds true: Let $(v_{q},\mathring{R}_{q})$ for some $q\in\mathbb{N}_{0}$ be an $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted solution to (3.4) satisfying
@@
$|\!|\!|\mathring{R}_{q}|\!|\!|_{L^{1},1}$ $\leqslant\frac{1}{48}c_{R}\delta_{q+2}\underline{e},$ (3.8f)
@@
$\frac{3}{4}\delta_{q+1}e(t)\leqslant e(t)-\mathbf{E}\|(v_{q}+z_{q})(t)\|_{L^{2}}^{2}\leqslant\frac{5}{4}\delta_{q+1}e(t),\quad$ $q\geqslant 1,$ (3.9)
$\mathbf{E}\|v_{0}(t)+z_{0}(t)\|^{2}_{L^{2}}\leqslant\frac{1}{2}e(t),\quad$ $q=0.$
@@
where $\bar{M}$ is a universal constant which will be fixed throughout the iteration.
@@
where we choose $a$ sufficiently large to absorb the constant in the last inequality. Taking expectation and using (3.7), we obtain
@@
where we used $6{\cdot}48{\cdot}(2\pi)^{3}c_{R}^{-1}L^{2}\leqslant\underline{e}$, $\delta_{1}=3rL^{2}$ and $\delta_{2}=\frac{1}{2}$. Hence, (3.8d), (3.8e) and (3.8f) are satisfied on the level $q=0$. Similarly, (3.9) also holds at the level $q=0$, since we have
@@
We will see in the following context, as long as the $q=0$ step in (3.9) holds ture, $\zeta_{0}(t)$ in (4.11) below is well-defined, thus the iteration can proceed.
@@ 3.2. Proof of main result
There are infinitely many such solutions by choosing different energies $e$.
@@
We may thus start the iteration Proposition 3.2 with the pair $(v_{0},\mathring{R}_{0})$ and obtain a sequence of solutions $(v_{q},\mathring{R}_{q})$. By (3.8c), (3.10) and interpolation we deduce for any $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}b^{5}},\frac{1}{21{\cdot}b^{4}}\}}\right)\in(0,\frac{2}{3}\beta)$, the following series is summable
@@
since $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{b+\sigma},\frac{8b-1}{9b}\}}$. Then we obtain $\lim_{q\to\infty}z_{q}=z$ in $L^{p}(\Omega;C({\mathbb{R}},C^{\vartheta}))$ for any $p\geqslant 1$. Furthermore, it follows from (3.8e) that $\lim_{q\to\infty}\mathring{R}_{q}=0$ in $L^{1}(\Omega;C({\mathbb{R}},C^{0}))$. Thus $v$ is an analytically weak solution to (3.2). Hence letting $u=v+z$ we obtain an $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted analytically weak solution to (1.1). Moreover, the estimate (3.11) for $u$ holds. Finally, (3.12) follows from (3.9), and choosing different energy functions, we can obtain infinitely many such solutions. This concludes the proof of Theorem 3.3. ∎
4. Proof of Proposition 3.2
The proof proceeds in several main steps which are similar as in many convex integration schemes. First of all, we start the construction by fixing the parameters in Section 4.1 and proceed with a mollification step in Section 4.2. Section 4.3 is devoted to introduce the new iteration $v_{q+1}$. This is the main part of the construction which differs in general convex integration schemes since we aim to construct global Hölder continuous flows. Here, we combine the construction methods of [BV19] and [HZZ22b] to obtain a new perturbation $w_{q+1}$. More precisely, we construct new amplitude functions $a_{(\xi)}$ similarly to [HZZ22b], meanwhile, we incorporate the solutions for transport equations into Beltrami waves to obtain an acceptable transport error similar to [BV19]. As pathwise positive lower bounds of the solutions to the transport equations are required, we use cutoff method to control the growth of noise and add pathwise estimtes of the velocity field $v_{q}$ in inductive iteration. Section 4.4 contains the inductive estimates of $v_{q+1}$, especially the pathwise and moment estimates. In Section 4.5, we show how to construct the new stress $\mathring{R}_{q+1}$. We establish the inductive pathwise and moment estimates on $\mathring{R}_{q+1}$ in Section 4.6. Finally, in Section 4.7 we show how the energy is controlled.

### v3

@@ 2.2. Probabilistic elements
For a given probability measure $\mathbf{P}$, we use $\mathbf{E}$ to denote the expectation under $\mathbf{P}$. Concerning the driving noise, we assume that $B$ is $\mathbb{R}^{3}$-valued two-sided $GG^{*}$-Wiener process with mean zero and divergence-free. This process is defined on some probability space $(\Omega,\mathcal{F},\mathbf{P})$, where $G$ is a Hilbert–Schmidt operator from $U$ to $L_{\sigma}^{2}$ for some Hilbert space $U$.
Given a Banach space $X=C(\mathbb{T}^{3},{\mathbb{R}}^{3})$ or $X=C^{\kappa}(\mathbb{T}^{3},{\mathbb{R}}^{3})$ for some $\kappa>0$. For $p\in[1,\infty)$ and $\delta\in(0,\frac{1}{2})$, we denote
$|\!|\!|u|\!|\!|^{p}_{X,p}:=\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{s\in[t,t+1]}\|u(s)\|^{p}_{X}\right],\quad|\!|\!|u|\!|\!|^{p}_{C_{t}^{\frac{1}{2}-\delta}X,p}:=\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|u\|^{p}_{C_{t}^{\frac{1}{2}-\delta}X}\right].$
The above norms denote function spaces of random variables on $\Omega$ taking values in $C({\mathbb{R}},X)$ and $C^{\frac{1}{2}-\delta}({\mathbb{R}},X)$, respectively, with bounds in $L^{p}(\Omega;{C(I,X))}$ and $L^{p}(\Omega;{C^{\frac{1}{2}-\delta}(I,X)})$ for bounded interval $I\subset{\mathbb{R}}$. Importantly, the bounds solely depend on the length of the interval $I$ and are independent of its location within ${\mathbb{R}}$. In addition, we denote the corresponding norms with $X$ replaced by $L^{1}$ and $H^{\frac{3}{2}+\kappa}$ for some $\kappa>0$.
@@ 3. Stochastic convex integration with pathwise estimates and results
In [CDZ24] and [HZZ22b], the authors presented a stochastic convex integration method to construct solutions on the entire timeline $\mathbb{R}$. They achieved this by introducing expectations to the iterative estimates in convex integration. In this section, we integrate their methods with pathwise estimates in the context of the stochastic Euler equations. Our objective is to construct Hölder continuous Euler flows, which are similar to the deterministic case as presented in [BV19]. In that case, the transport equations (see (4.3) below) are used to improve regularity. However, this method necessitates pathwise positive lower bounds for the solutions of the transport equations. Moment bounds during the inductive iteration become insufficient. Therefore, we introduce additional pathwise estimates in convex integration schemes. To control the growth of the noise, we employ the cut-off technique to the stochastic convolution. This enables the introduction of pathwise estimates during the inductive iteration. We combine this with moment bounds to derive desired uniform estimates. Compared to only using moment bounds in the inductive iteration in [CDZ24] and [HZZ22b], the advantage of our approach is that it only requires lower moments of the solutions, and avoids the requirements for higher moments as in [HZZ22b]. This can simplify the computations and enhance the regularity of the solutions.
We intend to develop an iteration procedure leading to the proof of Theorem 1.2. To this end, we decompose stochastic Euler system (1.1) into two parts, one is linear and involves the noise, whereas the second one is a random PDE. More precisely, we consider the stochastic linear equation
@@
where $B$ is $\mathbb{R}^{3}$-valued two-sided trace-class $GG^{*}$-Wiener process with divergence-free and mean zero, and $v$ solves the nonlinear equation
@@
Here, $z$ is divergence-free by the assumptions on the noise $B$ and we denote the pressure term associated with $v$ by $P$
Furthermore, using the factorization method, it is standard to obtain the regularity of $z$ on a given stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ being normal filtration given in [LR15, Section 2.1]. Specifically, the following result follows from [DPZ92, Theorem 5.16] together with the Kolmogorov continuity criterion.
@@
We recall that the unique stationary solution to (3.1) has an explicit form $z(t)=\int_{-\infty}^{t}e^{s-t}{\mathord{{\rm d}}}B_{s}$. According to [LR15, Section 2.1], the Wiener process $B$ can be written as $B=\sum_{k\in{\mathbb{N}}}\sqrt{c_{k}}\beta_{k}e_{k}$ for an orthonormal basis $\{e_{k}\}_{k\in{\mathbb{N}}}$ of $L^{2}_{\sigma}$ consisting of eigenvectors of $GG^{*}$ with corresponding eigenvalues $c_{k}$, and the coefficients satisfy $\sum_{k\in{\mathbb{N}}}c_{k}<\infty$. Here, $\{\beta_{k}\}_{k\in{\mathbb{N}}}$ denotes a sequence of mutually independent standard two-sided real-valued Brownian motions. Then it holds for $t\geqslant s$,
@@
By Kolmogorov’s continuity criterion and fundamental Sobolev embedding $H^{\frac{3}{2}+\kappa}\hookrightarrow C^{\kappa}$, we obtain for any $\delta\in(0,\frac{1}{2})$
@@ 3.1. Stochastic convex integration set-up and iterative proposition
Now, we apply the convex integration method to the nonlinear equation (3.2). The convex integration iteration is indexed by a parameter $q\in\mathbb{N}_{0}$. We consider an increasing sequence $\{\lambda_{q}\}_{q\in\mathbb{N}_{0}}$ which diverges to $\infty$, and a bounded sequence $\{\delta_{q}\}_{q\in\mathbb{N}}$ which decreases to 00. We choose $a,b\in\mathbb{N}$ sufficiently large and $\beta\in(0,1)$ sufficiently small and let
@@
More details on the choice of these parameters will be given in Section 4.1 below.
At each step $q$, a pair $(v_{q},\mathring{R}_{q})$ is constructed to solve the following system
@@
where $\chi_{q}$ is a non-increasing smooth function satisfying
$\chi_{q}(x)=\begin{cases}1,&x\in[0,\lambda_{q+2}^{\gamma}],\\ 0,&x\in(\lambda_{q+3}^{\frac{1}{4}\gamma},\infty).\end{cases}$
Moreover, we suppose $\chi_{q}$ with derivative bounded by $1$. This is achieved by choosing $a$ sufficiently large such that
$\frac{1}{\lambda_{q+3}^{\gamma/4}-\lambda_{q+2}^{\gamma}}\ll 1.$ (3.6)
In addition, the $\tilde{z}_{q}$ in (3.5) is given by $\tilde{z}_{q}=\mathbb{P}_{\leqslant f(q)}z$ with $f(q)=\lambda_{q+1}^{\gamma/8}$. Here $\mathbb{P}_{\leqslant f(q)}$ is the Fourier multiplier operator, which projects a function onto its Fourier frequencies $\leqslant f(q)$ in absolute value, and $\gamma$ is another small parameter which will be fixed in Section 4.1. $\mathring{R}_{q}$ on the right hand of (3.4) is a $3\times 3$ matrix which is trace-free and we put the trace part into the pressure. Keeping (3.3) in mind, by the Sobolev embedding $\|f\|_{L^{\infty}}\lesssim\|f\|_{H^{\frac{3}{2}+\kappa}}$ for $\kappa>0$, we observe that for any $p\geqslant 2$ and $\delta\in(0,\frac{1}{2})$
@@
The above estimates (3.7) will be frequently used in the sequel.
Remark 3.2 (On the choice of cut-off).
In order to apply the stationary phase lemma (see Lemma 5.3 below) to derive the desired bound of $\mathring{R}_{q+1}$, it is necessary to have a $pathwise$ positive lower bound for the solution $\Phi_{k,j}$ of equation (4.3) below (as shown in (4.4b)). This requires a pathwise bound of $C^{1}_{x}$-norm of $z_{q}$ part from the noise. Consequently, we truncate the $C^{1}_{x}$-norm of $\tilde{z}_{q}$ by a cut-off function depending on $q$. Although the pathwise $C^{1}_{x}$-norm for $z_{q}$ may blow up, we choose the parameter $\lambda_{q+3}^{\gamma/4}$ in the definition of $\chi_{q}$ to ensure it does not increase too fast. On the other hand, our technique requires absorbing the difference of $z_{q+1}$ and $z_{q}$, i.e. $|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},r}$ into the small parameter $\delta_{q+3}$ to derive the inductive estimates for $\mathring{R}_{q+1}$ (see (3.8e) and (3.8f) below). Therefore, the choice of parameter $\lambda_{q+2}^{\gamma}$ is to bound $|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},r}$ by $\lambda_{q+2}^{-\gamma}$, which can be absorbed into $\delta_{q+3}$ by choosing suitable parameters as stated in Section 4.1 below.
Under the above assumptions, the key result is the following iterative proposition, and the proof of this result is presented in Section 4 and Section 5 below.
@@
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$ and let $r>1$ be fixed. Given a smooth function $e:\mathbb{R}\to(0,\infty)$ such that $\bar{e}\geqslant e(t)\geqslant\underline{e}\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}L^{2}$ with $\|e^{\prime}\|_{C^{0}}\leqslant\tilde{e}$ for some constants $\overline{e},\underline{e},\tilde{e}>0$, where $L$ is from Proposition 3.1 and $c_{R}>0$ is a small universal constant (see (4.8) below). There exists a choice of parameters $a,b,\beta$ and $\gamma$ such that the following holds true: Let $(v_{q},\mathring{R}_{q})$ for some $q\in\mathbb{N}_{0}$ be an $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted solution to (3.4) satisfying
@@
$|\!|\!|\mathring{R}_{q}|\!|\!|_{L^{1},1}$ $\leqslant\frac{1}{48}c_{R}\delta_{q+2}\underline{e}.$ (3.8f)
@@
$\frac{3}{4}\delta_{q+1}e(t)\leqslant e(t)-\mathbf{E}\|(v_{q}+z_{q})(t)\|_{L^{2}}^{2}\leqslant\frac{5}{4}\delta_{q+1}e(t),$ $q\geqslant 1,$ (3.9)
$\mathbf{E}\|v_{0}(t)+z_{0}(t)\|^{2}_{L^{2}}\leqslant\frac{1}{2}e(t),$ $q=0.$
@@
where $\bar{M}$ is a universal constant that will be fixed throughout the iteration.
@@
where we chose $a$ sufficiently large to absorb the constant in the last inequality. Taking expectation and using (3.7), we obtain
@@
where we used $6{\cdot}48{\cdot}(2\pi)^{3}c_{R}^{-1}L^{2}\leqslant\underline{e}$, $\delta_{1}=3rL^{2}$ and $\delta_{2}=\frac{1}{2}$. Hence, (3.8d), (3.8e) and (3.8f) are satisfied at the level $q=0$. Similarly, (3.9) also holds at the level $q=0$, since we have
@@
We will see in the following context, as long as the $q=0$ step in (3.9) holds, $\zeta_{0}(t)$ in (4.7) below is well-defined, thus the iteration can proceed.
@@ 3.2. Proof of main result
There are infinitely many such solutions by choosing different energy functions $e$.
@@
We start the iteration Proposition 3.3 with the pair $(v_{0},\mathring{R}_{0})$ and obtain a sequence of solutions $(v_{q},\mathring{R}_{q})$. By (3.8c), (3.10) and interpolation we deduce for any $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}b^{5}},\frac{1}{21{\cdot}b^{4}}\}}\right)\in(0,\frac{2}{3}\beta)$, the following series is summable
@@
where we used $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{b+\sigma},\frac{8b-1}{9b}\}}$ in the last inequality. Then, for any $p\geqslant 1$, we have $\lim_{q\to\infty}z_{q}=z$ in $L^{p}(\Omega;C({\mathbb{R}},C^{\vartheta}))$. Furthermore, it follows from (3.8e) that $\lim_{q\to\infty}\mathring{R}_{q}=0$ in $L^{1}(\Omega;C({\mathbb{R}},C^{0}))$. Thus, $v$ is an analytically weak solution to (3.2). Letting $u=v+z$ we obtain an $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted analytically weak solution to (1.1) and the estimate (3.11) for $u$ holds. Finally, (3.12) follows from (3.9). By choosing different energy functions, we can obtain infinitely many such solutions. This concludes the proof of Theorem 3.4. ∎
Remark 3.5.
It appears challenging to specify the energy pathwise, as done in [HZZ23b]. Indeed, to obtain the pathwise estimates for energy, our technique requires absorbing the pathwise $C^{0}_{x}$-norms of $\mathring{R}_{q}$ and $z_{q+1}-z_{q}$ into the small parameter $\delta_{q+2}$, as demonstrated in the proof of Proposition 5.4 below. Instead of using stopping time as in [HZZ23b] to control the noise, we utilize a cut-off function $\chi_{q}$ here, the support of which also increases with $q$. This makes it not easy to absorb the above two terms into $\delta_{q+2}$.
4. Convex integration scheme
In this section, we aim to construct a pair $(v_{q+1},\mathring{R}_{q+1})$ to solve the system (3.4), given $(v_{q},\mathring{R}_{q})$ as in the statement of Proposition 3.3. First of all, we fix the parameters used during the construction in Section 4.1 and continue with a mollification step in Section 4.2. Section 4.3 is devoted to introducing the new perturbation $w_{q+1}$. We follow the construction method in [HZZ22b] and [BV19] to achieve this. More precisely, we construct new amplitude functions $a_{(\xi)}$ similarly to [HZZ22b], meanwhile, we incorporate the solutions for transport equations into Beltrami waves to obtain an acceptable transport error as in [BV19]. As pathwise positive lower bounds of the solutions to the transport equations are required, we use the cut-off technique to control the growth of noise and add pathwise estimates of the velocity field $v_{q}$ in inductive iteration. In Section 4.4, we show how to construct the new stress $\mathring{R}_{q+1}$.